OpenStudy (anonymous):

how do I solve compound inequalities -8k+7<-7k+2 or -6+2k greater than =to 3k-6

OpenStudy (aroub):

First solve each one alone

OpenStudy (aroub):

-8k+7<-7k+2 -8k+7k<-7+2 -k<-5 k>5

OpenStudy (anonymous):

do I combine the like terms?

OpenStudy (aroub):

-6+2k>=3k-6 -3k+2k>=-6+6 -k>=0 k>0

OpenStudy (aroub):

So, k>5 or k>0

OpenStudy (aroub):

Yes! You add like terms

OpenStudy (anonymous):

thank you I got it now we did not do any in class

OpenStudy (aroub):

Oh! Maybe in class you'll understand it more :) Good luck!